OpenStudy (anonymous):

Simplify: -5(sqaure root symbol)112

OpenStudy (anonymous):

\[-5\sqrt{112}\]

OpenStudy (jdoe0001):

\(\bf -5\sqrt{112} \\ \quad \\ {\color{brown}{ 112\to 2\cdot 2\cdot 2\cdot 2\cdot 7\to 2^2\cdot 2^2\cdot 7\to (2^2)^2\cdot 7}}\qquad thus \\ \quad \\ -5\sqrt{112}\implies -5\sqrt{(2^2)^2\cdot 7}\implies -5\sqrt{(2^2)^2}\cdot \sqrt{7}\implies ?\)

OpenStudy (anonymous):

how do i solve that

OpenStudy (jdoe0001):

well... what's \(\Large \sqrt[{\color{red}{ 2}}]{x^{\color{red}{ 2}}}\implies ?\)

OpenStudy (anonymous):

Use a factor tree, go on khan acdemy, im sure its there :P

OpenStudy (anonymous):

i know how to do a factor tree i just get confused after that

OpenStudy (jdoe0001):

so.... when taking roots..... what \(\Large \sqrt[{\color{red}{ 2}}]{x^{\color{red}{ 2}}}\implies ?\) would be? notice, the radicand exponent "matches" the root

OpenStudy (anonymous):

so itd be 16?

OpenStudy (jdoe0001):

well \(\bf (2^2)^2\) is 16 yes.. but once "out of the radical" is not